• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Tyler Hague

    Log entry time 07:49:29 on December 7,2014

    Entry number 404648

    keyword=Target and strip charts

    Target and strip charts. Target is stable at 19K. JT valve issues from earlier in the night have not recurred.

    I now believe that something causes the IOC to lose communication with the JT valve when the heater power gets too low. Since we have adjusted the valve to keep the heater from running at too low a power, the problem has not come back.

    The heater currently has the max power set at 400W. We had the heater running at a higher reserve power for when we recieved 20 uA beam. This was set after talking with Silviu. We are no longer keeping the heater running that high because we have only been getting 5 uA beam.


    FIGURE 1